How much paint is needed for the inside and outside painting of 300 buckets having a shape of a frustum of a cone, if the diameter of the bases of the buckets are 32 in. and 38 in. and its height is 45 in. ?Consider 5% waste and a 1 liter per 500 sq. in for paint consumption.
Answer by mananth(16946)

area, not including the top and bottom circles, is
A = pi(R_1+R_2)s


= pi(R_1+R_2)sqrt((R_1-R_2)^2+h^2).
= 3.142*(70)*sqrt(6^2+45^2)
=9984.88
Area of Base =pi*r2^2
= 3.142 * 38^2
=4537
Total area = 9984.88+4537
=14521.88 in^2
both sides
14521.88 in^2 *2
29043.76 in^2
1 litre = 500 in^2
Number of litre = 58 litres
Add 5% waste
=61 litres